# Spectroscopy using a visible photonic lantern at the Subaru telescope: Laboratory characterization and first on-sky demonstration on Ikiiki ( Leo) and ‘Aua ( Ori)

### v1

@@ Abstract
Context. Photonic lanterns are waveguide devices enabling high throughput single mode spectroscopy and high angular resolution.
Aims. We aim to present the first on-sky demonstration of a photonic lantern (PL) operating in visible light, to measure its throughput and assess its potential for high-resolution spectroscopy of compact objects.
Methods. We used the SCExAO instrument (a double stage extreme AO system installed at the Subaru telescope) and FIRST mid-resolution spectrograph (R 3000) to test the visible capabilities of the PL on internal source and on-sky observations.
Results. The best averaged coupling efficiency over the PL field of view was measured at with a peak at . We also investigate the relationship between coupling efficiency and the Strehl ratio for a PL, comparing them with those of a single-mode fiber (SMF). Findings show that in the AO regime, a PL offers better coupling efficiency performance than a SMF, especially in the presence of low spatial frequency aberrations. We observed Ikiiki ( Leo - = 1.37) and ‘Aua ( Ori - = -1.17) at a frame rate of 200 Hz. Under median seeing conditions (about 1 arcsec measured in H-band) and large tip/tilt residuals (over 20 mas), we estimated an average light coupling efficiency of , with a maximum of at 680 nm. We were able to reconstruct both star’s spectra, containing various absorption lines.
Conclusions. The successful demonstration of this device opens new possibilities in terms of high throughput single-mode fiber-fed spectroscopy in the Visible. The demonstrated on-sky coupling efficiency performance would not have been achievable with a single SMF injection setup under similar conditions, partly because the residual tip/tilt alone exceeded the field of view of a visible SMF (18 mas at 700 nm). Thus emphasizing the enhanced resilience of PL technology to such atmospheric disturbances. The additional capabilities in high angular resolution are also promising but still have to be demonstrated in a forthcoming investigation.
@@ 1 Introduction
Spectroscopy at high throughput and high resolution plays a key role in the study of the universe, whether it is used for exoplanet detection/characterization, stellar physics or cosmology. High dispersion spectroscopic (HDS) observations of star-planet systems have recently been used to isolate giant exoplanet signals via Doppler shifts, enabling measurements of orbital velocities, masses, and molecular features (e.g. Snellen et al., 2010; Brogi et al., 2012; Currie et al., 2023a). This HDS analysis approach, coupled with high contrast imaging, is also expected to enable rocky planet characterization with next-generation telescopes (Snellen et al., 2015). High spectral resolution observations of stars have enabled a broad range of science, from measurements of circumstellar disk and wind kinematics (e.g. Pontoppidan et al., 2011) and ices (e.g. Boogert et al., 2002), to tests of general relativity via radial velocity measurements of stars in the galactic center (e.g. Do et al., 2019). High-throughput, high-resolution spectrographs are also crucial for a wide range of extragalactic faint source applications, from transient follow-up (e.g. Patat et al., 2007), to characterizing stellar populations (e.g. Reddy et al., 2022), to intermediate mass black hole searches (e.g. Casares & Jonker, 2014, and references therein).
@@
These broad and impactful applications motivate the development of technologies that can maximize throughput while preserving high spectral resolution and quality. Compared to traditional multi-mode or slit-based spectroscopy, single-mode fiber (SMF) fed spectroscopy offers numerous advantages for astronomical observations, ranging from improved spectral quality and resolution to increased instrument compactness and stability. The beam coming out of an SMF is indeed highly stable in shape, since it is determined by the fundamental mode of the fiber, ensuring a stable spectral response of the spectrograph. The weakness of single-mode fibers lies in the challenge of light injection, requiring very fine alignment of the source to the fiber core and high Strehl ratio - which are difficult to obtain especially in visible light. Light from a telescope can be efficiently injected into an SMF with the help of extreme adaptive optics (ExAO) (Jovanovic et al., 2017). ExAO systems are designed to measure and correct wavefront aberrations induced by atmospheric turbulence, enabling Strehl ratios up to in the -band.
@@
Multiple post-ExAO SMF-based instruments have been or will be commissioned on major observatories: Subaru/REACH (Kotani et al., 2020), Keck/KPIC (Delorme et al., 2021), Subaru/GLINT (Martinod et al., 2021), VLT/HiRISE (Vigan et al., 2024) and Subaru/FIRST (Vievard et al., 2023a). While some of them differ by their wavelength operation (Subaru/FIRST in the visible, the rest in the Near-IR) all share the same struggle of coupling efficiency/throughput limitations.
@@
A solution that combines high coupling efficiency with the stability and compactness advantages of SMF-based spectroscopy was proposed by Leon-Saval et al. (2005). It consists of using a device transitioning from a multi-mode (MM) input to multiple single-mode (SM) outputs. The MM input ensures high coupling efficiency, the slow transition between MM and SM ensures high throughput and the SM outputs allows use of SMF-fed spectroscopy. Such devices are called Photonic Lanterns (PL). The interest of PLs also lies in their sensitivity to the input scene and wavefront: the flux distribution among the PL outputs depends both on the source position/shape (Kim et al., 2022) and on the wavefront input (Norris et al., 2020). The corresponding information allows for retrieval of the geometry/position of the observed object, as well as wavefront sensing and optimization of injection efficiency into one specific output SMF if desired (Norris et al., 2022). These features are not possible with SMFs hence PLs can potentially provide access to a wider range of science applications.
@@
The characteristics described above make PLs an attractive technology for high-contrast imaging science. One particular example is spatially- and spectrally-resolved searches for accreting protoplanets orbiting nearby young stars. Conventional imaging of these targets is typically limited to mas angular separations (Keppler et al., 2018; Haffert et al., 2019; Currie et al., 2022), corresponding to orbital separation of AU for targets at 150 pc. Simulations of both spectroastrometry and interferometric imaging with PLs have shown that they can be used to resolve few AU orbits around nearby young stars (Xin et al., 2022; Kim et al., 2022; Levinstein et al., 2023; Kim et al., 2024) - a region where the majority of known mature giant planets reside (Fernandes et al., 2019). Coupling the SMF outputs to a medium resolution spectrograph also allows for spectral-differential measurements that can isolate accretion-tracing hydrogen lines (Kim et al., 2022; Levinstein et al., 2023). These excesses, along with spectral slopes, can distinguish between scattered light by circumstellar disk material and self-luminous planets (e.g. Haffert et al., 2019; Currie et al., 2022; Wagner et al., 2023), which is crucial for bona-fide protoplanet detections. Beyond protoplanets, PLs could pay dividends for a broad range of science cases requiring high angular resolution observations, where interferometry is often utilized, including studies of close-in spirals in protoplanetary disks (e.g. Sallum et al., 2019), inner regions of AGN tori (e.g. Ford et al., 2014), and Solar System object surfaces (e.g. Conrad et al., 2015).
@@
We recently acquired a PL with 19 SMF outputs, designed for visible wavelengths. The device was installed at the 8.2 m Subaru Telescope as a sub-module of the Subaru Extreme Adaptive Optics (SCExAO) platform to explore the possibilities of using PLs for high throughput and high resolution spectroscopy. Stellar light can be injected into the PL in a pupil plane or in a focal plane. This paper focuses on the latter approach, though the former is also a possibility. The 19 PL outputs feed the Subaru/FIRST mid-resolution spectrograph (R) optimized for wavelengths from 600 to 800 nm (Vievard et al., 2023b). A single-mode fiber is also installed in parallel and feeds the same spectrograph for relative comparison between the two methods.
In section 2 we provide a comprehensive description of the PL device, and its integration on SCExAO. Section 3 presents in-lab characterization of light injection within the PL. Finally, section 4 presents the first demonstration of spectroscopy in the Visible wavelengths, performed using an ExAO-fed PL. We observe and reconstruct the spectrum of two targets: Ikiiki 111This paper refers to the main targets per their Hawaiian names, in honor of Mauna Kea, the mountain from which we were most fortunate to perform our observations, and with acknowledgment of the indigenous hawaiian communities who have had and continue to have Kuleana (translated into English as ”responsibility”) to this land. ( Leo) and ‘Aua ( Ori).
@@ 2.1 The 19 pigtailed outputs Photonic Lantern
The PL used in this experiment has one MM input and 19 SM outputs. The SM outputs of a PL come in various configurations (Leon-Saval et al., 2010). Our lantern was constructed from separate SMFs where the outputs are separate pigtails (see drawing in Figure 1). An adiabatic taper allows for an efficient and low-loss mode conversion (throughput greater than 90%; Birks et al., 2015). This is accomplished by placing the multiple SM fibers into a low refractive-index capillary (having an index less than that of the SM fiber cladding, which in turn is less than that of the SM fiber core). This assembly is tapered down, such that the cores of the SM fibers disappear, their cladding becomes the new core of a MM region and the capillary becomes the MM region’s cladding. This smooth transition is crucial since abrupt changes in the waveguide geometry can lead to internal reflections and mode mismatch which would result in decreasing the device throughput. In the transition region, light propagating in the modes supported by the input waveguide section is redistributed across the modes of the output - this process is reciprocal. As long as the number of output SM fibers is equal to or exceeds the number of unique modes supported in the input MM region, this process has very high efficiency. The PL used in this experiment was designed to work at the lower end range of the FIRST wavelength coverage (around 650 nm). This design choice ensures that, at the shortest wavelengths, the number of modes does not increase significantly, which would otherwise contribute to transmission losses in the PL device. The 19 SMFs are then spliced to a linear V-groove allowing a spectrograph to cross-disperse the light.
@@ 2.2 Instrument setup: integration on SCExAO
The Subaru Coronagraphic Extreme Adaptive Optics (SCExAO - Jovanovic et al., 2015) instrument is located on the Nasmyth IR plateform at the 8.2-meter Subaru Telescope. SCExAO receives light from the Subaru Telescope AO facility instrument, AO188 (Minowa et al., 2010). AO188 performs a first stage of AO correction using a curvature wavefront sensor (CWFS) controlling a 188-element deformable mirror (DM). SCExAO performs a second stage of AO correction, in order to reach the ExAO regime (with Strehl ratio in the 0.8–0.9 range in H-band) therefore enabling high contrast imaging at high angular resolution and the detection of faint planets and brown dwarf companions (Kuzuhara et al., 2022; Currie et al., 2023b; Tobin et al., 2024). To achieve ExAO corrections, a pyramid wavefront sensor (PyWFS - Lozi et al., 2019) operating in the nm to nm wavelength range drives a 2040-actuator MEMS DM.
Figure 3 shows the optical configuration bringing the light from the telescope to the PL and the spectrograph. Light coming from AO188 is collimated by an off-axis parabola (OAP). A fiber carrying light from a Supercontinuum laser (hereafter named SuperK) can be inserted at the focal point for calibration/alignment purposes. The collimated light is then reflected by the 2040-actuator DM toward a pupil mask - mimicking the Subaru Telescope central obstruction and spiders for off-sky calibration. A dichroic mirror then splits the light, with nearIR light ( nm) transmitted to the IR modules, and visible light ( nm) reflected to a periscope. The periscope allows for the visible instruments to be mounted on an optical bench above the bench hosting the IR modules.
After exiting the periscope, visible light is first split by a dichroic filter: the reflected beam (from nm to nm) is sent to the pyramid WFS and the transmitted beam (¡ nm) is used by the visible modules of SCExAO. Before picking off the light for the PL, a R90/T10 filter can be inserted to either send the light to the FIRST spectro-interferometer (Vievard et al., 2023a), or send light from the wavelength calibration unit (WCU) to the PL. The WCU is a fiber connected to a Neon lamp (AvaLight-CAL-Neon- Mini), used for the PL spectrograph wavelength calibration. A beamsplitter (BS) wheel allows reflecting the beam toward the PL, or transmitting the beam to the VAMPIRES dual-camera imager. The different slots of the BS wheel are currently occupied by a non-polarizing BS cube (NPBS), a polarizing BS cube, a nm short pass filter, a nm short pass filter, or an open slot. The following sub-sections present the injection module of the PL and the spectrograph.
@@ 2.2.1 The injection module
Figure 3 presents the optical configuration to obtain a focal ratio of about 4. After the BS, the beam is collimated by a f400 mm doublet (named L1 on Figure 3). Two additional doublets of focal length 400 mm (L2) and 50 mm (L3) focus the beam onto the PL MM fiber input (see Figure 4 Left). A folding mirror located between L2 and L3 folds the beam for compactness and alignment requirements. Doublets are achromatic, making them well suited for injecting into the MM end of the PL. As shown on Figure 4 (Left), the PL is mounted on a bracket composed by 2 ports. In the second port, an SMF is installed for comparison between the two devices. The bracket is mounted on two remotely controlled translation stages (Zaber Technologies) for motion in the plane perpendicular to the optical axis (X and Y axes in Figure 4 (Left)), and on a conex stage for focus motion (Z axis in Figure 4 (Left)). They allow the PL input to be scanned across the focal plane of the injection lens, in order to optimize the light coupling into the PL.

### v3

# Spectroscopy using a visible photonic lantern at the Subaru Telescope: Laboratory characterization and the first on-sky demonstration on Ikiiki ( Leo) and ‘Aua ( Ori)

@@ Abstract
Context. Photonic lanterns (PLs) are waveguide devices enabling high-throughput single-mode spectroscopy and high angular resolution.
Aims. We aim to present the first on-sky demonstration of a PL operating in visible light, to measure its throughput and assess its potential for high-resolution spectroscopy of compact objects.
Methods. We used the SCExAO instrument (a double-stage extreme adaptive optics system installed at the Subaru Telescope) and FIRST mid-resolution spectrograph (R 3000) to test the visible capabilities of the PL on internal source and on-sky observations.
Results. The best averaged coupling efficiency over the PL field of view was measured at , with a peak at . We also investigated the relationship between coupling efficiency and the Strehl ratio for a PL, comparing them with those of a single-mode fiber (SMF). Findings show that in the adaptive optics regime a PL offers a better coupling efficiency performance than an SMF, especially in the presence of low-spatial-frequency aberrations. We observed Ikiiki ( Leo - = 1.37) and ‘Aua ( Ori - = -1.17) at a frame rate of 200 Hz. Under median seeing conditions (about 1 arcsec measured in the H band) and large tip or tilt residuals (over 20 mas), we estimated an average light coupling efficiency of , with a maximum of at 680 nm. We were able to reconstruct both star’s spectra, containing various absorption lines.
Conclusions. The successful demonstration of this device opens new possibilities in terms of high-throughput single-mode fiber-fed spectroscopy in the visible. The demonstrated on-sky coupling efficiency performance would not have been achievable with a single SMF injection setup under similar conditions, partly because the residual tip or tilt alone exceeded the field of view of a visible SMF (18 mas at 700 nm). This emphasizes the enhanced resilience of PL technology to such atmospheric disturbances. The additional capabilities in high angular resolution are also promising but still have to be demonstrated in a forthcoming investigation.
@@ 1 Introduction
Spectroscopy at high throughput and high resolution plays a key role in the study of the Universe, whether it is used for exoplanet detection or characterization, stellar physics, or cosmology. High-dispersion spectroscopic (HDS) observations of star-planet systems have recently been used to isolate giant exoplanet signals via Doppler shifts, enabling measurements of orbital velocities, masses, and molecular features (e.g., Snellen et al., 2010; Brogi et al., 2012; Currie et al., 2023a). This HDS analysis approach, coupled with high contrast imaging, is also expected to enable rocky planet characterization with next-generation telescopes (Snellen et al., 2015). High-spectral-resolution observations of stars have enabled a broad range of science, from measurements of circumstellar disk and wind kinematics (e.g., Pontoppidan et al., 2011) and ices (e.g., Boogert et al., 2002), to tests of general relativity via radial velocity measurements of stars in the galactic center (e.g., Do et al., 2019). High-throughput, high-resolution spectrographs are also crucial for a wide range of extragalactic faint source applications, from transient follow-up (e.g., Patat et al., 2007), to characterizing stellar populations (e.g., Reddy et al., 2022), to searches for intermediate-mass black holes (e.g., Casares & Jonker, 2014, and references therein).
@@
These broad and impactful applications motivate the development of technologies that can maximize throughput, while preserving high spectral resolution and quality. Compared to traditional multimode (MM) or slit-based spectroscopy, single-mode fiber (SMF) fed spectroscopy offers numerous advantages for astronomical observations, ranging from improved spectral quality and resolution to increased instrument compactness and stability. The beam coming out of an SMF is indeed highly stable in shape, since it is determined by the fundamental mode of the fiber, ensuring a stable spectral response of the spectrograph. The weakness of SMFs lies in the challenge of light injection, requiring very fine alignment of the source to the fiber core and a high Strehl ratio, which are difficult to obtain, especially in visible light. Light from a telescope can be efficiently injected into an SMF with the help of extreme adaptive optics (ExAO) (Jovanovic et al., 2017). The ExAO systems are designed to measure and correct wavefront aberrations induced by atmospheric turbulence, enabling Strehl ratios up to in the band.
@@
Multiple post-ExAO SMF-based instruments have been or will be commissioned on major observatories: Subaru/REACH (Kotani et al., 2020), Keck/KPIC (Delorme et al., 2021), Subaru/GLINT (Martinod et al., 2021), VLT/HiRISE (Vigan et al., 2024), and Subaru/FIRST (Vievard et al., 2023a). While some of them differ by their wavelength operation (Subaru/FIRST is in the visible, the rest are in the near-infrared), all of them share the same struggle of coupling efficiency and throughput limitations.
@@
A solution that combines high coupling efficiency with the stability and compactness advantages of SMF-based spectroscopy was proposed by Leon-Saval et al. (2005). It consists of using a device transitioning from an MM input to multiple single-mode (SM) outputs. The MM input ensures high coupling efficiency, the slow transition between MM and SM ensures high throughput, and the SM outputs enable the use of SMF-fed spectroscopy. Such devices are called photonic lanterns (PLs). The interest in PLs also lies in their sensitivity to the input scene and wavefront: the flux distribution among the PL outputs depends both on the source position and shape (Kim et al., 2022) and on the wavefront input (Norris et al., 2020). The corresponding information allows for retrieval of the geometry and position of the observed object, as well as wavefront sensing and optimization of injection efficiency into one specific output SMF if desired (Norris et al., 2022). These features are not possible with SMFs; hence, PLs can potentially provide access to a wider range of science applications.
@@
The characteristics described above make PLs an attractive technology for high-contrast imaging science. One particular example is spatially and spectrally resolved searches for accreting protoplanets orbiting nearby young stars. Conventional imaging of these targets is typically limited to mas angular separations (Keppler et al., 2018; Haffert et al., 2019; Currie et al., 2022), corresponding to an orbital separation of AU for targets at 150 pc. Simulations of both spectroastrometry and interferometric imaging with PLs have shown that they can be used to resolve a few AU orbits around nearby young stars (Xin et al., 2022; Kim et al., 2022; Levinstein et al., 2023; Kim et al., 2024) – a region where the majority of known mature giant planets reside (Fernandes et al., 2019). Coupling the SMF outputs to a medium-resolution spectrograph also allows for spectral-differential measurements that can isolate accretion-tracing hydrogen lines (Kim et al., 2022; Levinstein et al., 2023). These excesses, along with spectral slopes, can distinguish between scattered light from circumstellar disk material and self-luminous planets (e.g., Haffert et al., 2019; Currie et al., 2022; Wagner et al., 2023), which is crucial for bona fide protoplanet detections. Beyond protoplanets, PLs could pay dividends for a broad range of science cases requiring high-angular-resolution observations, where interferometry is often utilized, including studies of close-in spirals in protoplanetary disks (e.g., Sallum et al., 2019), inner regions of active galactic nucleus tori (e.g., Ford et al., 2014), and Solar System object surfaces (e.g., Conrad et al., 2015).
@@
We recently acquired a PL with 19 SMF outputs, designed for visible wavelengths. The device was installed at the 8.2 m Subaru Telescope as a sub-module of the Subaru Extreme Adaptive Optics (SCExAO) platform to explore the possibilities of using PLs for high-throughput and high-resolution spectroscopy. Stellar light can be injected into the PL in a pupil plane or in a focal plane. This paper focuses on the latter approach, though the former is also a possibility. The 19 PL outputs feed the Subaru/FIRST mid-resolution spectrograph (R) optimized for wavelengths from 600 to 800 nm (Vievard et al., 2023b). An SMF is also installed in parallel and feeds the same spectrograph for a comparison between the two methods.
In Sect. 2, we provide a comprehensive description of the PL device, and its integration on SCExAO. Section 3 presents an in-lab characterization of light injection within the PL. Finally, Sect. 4 presents the first demonstration of spectroscopy in visible wavelengths, performed using an ExAO-fed PL. We observe and reconstruct the spectrum of two targets: Ikiiki 111This paper refers to the main targets per their Hawaiian names, in honor of Mauna Kea, the mountain from which we were most fortunate to perform our observations, and with acknowledgment of the indigenous hawaiian communities who have had and continue to have Kuleana (translated into English as ”responsibility”) to this land. ( Leo) and ‘Aua ( Ori).
@@ 2.1 The 19 pigtailed outputs Photonic Lantern
The PL used in this experiment has 1 MM input and 19 SM outputs. The SM outputs of a PL come in various configurations (Leon-Saval et al., 2010). Our lantern was constructed from separate SMFs where the outputs are separate pigtails (see the drawing in Fig. 1). An adiabatic taper allows for an efficient and low-loss mode conversion (throughput greater than 90%; Birks et al., 2015). This was accomplished by placing the multiple SM fibers into a low refractive-index capillary (that has an index less than that of the SM fiber cladding, which in turn is less than that of the SM fiber core). This assembly is tapered down, such that the cores of the SM fibers disappear, their cladding becomes the new core of a MM region, and the capillary becomes the MM region’s cladding. This smooth transition is crucial, since abrupt changes in the waveguide geometry can lead to internal reflections and mode mismatch, which would result in decreasing the device throughput. In the transition region, light propagating in the modes supported by the input waveguide section is redistributed across the modes of the output - this process is reciprocal. As long as the number of output SM fibers is equal to or exceeds the number of unique modes supported in the input MM region, this process has very high efficiency. The PL used in this experiment was designed to work at the lower end of the FIRST wavelength coverage (around 650 nm). This design choice ensures that, at the shortest wavelengths, the number of modes does not increase significantly, which would otherwise contribute to transmission losses in the PL device. The 19 SMFs were then spliced to a linear V-groove (VG), allowing a spectrograph to cross-disperse the light.
@@ 2.2 Instrument setup: Integration on SCExAO
The Subaru Coronagraphic Extreme Adaptive Optics (SCExAO - Jovanovic et al., 2015) instrument is located on the Nasmyth infrared (IR) platform at the 8.2-meter Subaru Telescope. SCExAO receives light from the Subaru Telescope AO facility instrument, AO188 (Minowa et al., 2010). AO188 performs a first stage of AO correction using a curvature wavefront sensor (CWFS) controlling a 188-element deformable mirror (DM). SCExAO performs a second stage of AO correction, in order to reach the ExAO regime (with a Strehl ratio in the 0.8–0.9 range in the H band), enabling high-contrast imaging at high angular resolution and the detection of faint planets and brown dwarf companions (Kuzuhara et al., 2022; Currie et al., 2023b; Tobin et al., 2024). To achieve ExAO corrections, a pyramid wavefront sensor (PyWFS - Lozi et al., 2019) operating in the nm to nm wavelength range drives a 2040-actuator MEMS DM.
Figure 3 shows the optical configuration bringing the light from the telescope to the PL and the spectrograph. Light coming from AO188 is collimated by an off-axis parabola (OAP). A fiber carrying light from a Supercontinuum laser (hereafter named SuperK) can be inserted at the focal point for calibration or alignment purposes. The collimated light is then reflected by the 2040-actuator DM toward a pupil mask - mimicking the Subaru Telescope central obstruction and spiders for off-sky calibration. A dichroic mirror then splits the light, with near-infrared light ( nm) transmitted to the IR modules, and visible light ( nm) reflected to a periscope. The periscope allows for the visible instruments to be mounted on an optical bench above the bench hosting the IR modules.
After exiting the periscope, visible light is first split by a dichroic filter: the reflected beam (from nm to nm) is sent to the pyramid WFS and the transmitted beam (¡ nm) is used by the visible modules of SCExAO. Before picking off the light for the PL, a R90/T10 filter can be inserted to either send the light to the FIRST spectro-interferometer (Vievard et al., 2023a), or send light from the wavelength calibration unit (WCU) to the PL. The WCU is a fiber connected to a neon lamp (AvaLight-CAL-Neon- Mini), used for the PL spectrograph wavelength calibration. A beamsplitter (BS) wheel allows the beam to be reflected toward the PL or transmitted to the VAMPIRES dual-camera imager. The different slots of the BS wheel are currently occupied by a non-polarizing BS cube (NPBS), a polarizing BS cube, a nm short pass filter, a nm short pass filter, or an open slot. The following subsections present the injection module of the PL and the spectrograph.
@@ 2.2.1 The injection module
Figure 3 presents the optical configuration to obtain a focal ratio of about 4. After the BS, the beam is collimated by a f400 mm doublet (named L1 in Fig. 3). Two additional doublets of focal length 400 mm (L2) and 50 mm (L3) focus the beam onto the PL MM fiber input (see Fig. 4 Left). A folding mirror located between L2 and L3 folds the beam for compactness and alignment requirements. Doublets are achromatic, making them well suited for injecting into the MM end of the PL. As is shown in Fig. 4 (Left), the PL is mounted on a bracket composed of two ports. In the second port, an SMF is installed for comparison between the two devices. The bracket is mounted on two remotely controlled translation stages (Zaber Technologies) for motion in the plane perpendicular to the optical axis (X and Y axes in Fig. 4 (Left)), and on a conex stage for focus motion (Z axis in Fig. 4 (Left)). They allow the PL input to be scanned across the focal plane of the injection lens, in order to optimize the light coupling into the PL.